OpenStudy (anonymous):

How do I change this point slope equation into standard? y - 3/5 = -4/5 (x + 6)

OpenStudy (aem9296):

Standard form is written as Ax+By=C So, when looking at your equation you first multiply out the -4/5 to the numbers in the parenthesis \[y - \frac{ 3 }{ 5 }=-\frac{ 4 }{ 5 }x - \frac{ 24 }{ 5 }\] Then you add3/5 and 4/5x to both sides. \[y+\frac{ 4 }{ 5}x=-\frac{ 21 }{ 5}\] Yet, in standard form the variables have to be whole numbers. So, multiply both sides by 5. \[5y+4x=-21\]

OpenStudy (anonymous):

How did you get rid of the 3/5

OpenStudy (aem9296):

you add it to both sides

OpenStudy (anonymous):

OHHHHH so you add it onto the -21/5

OpenStudy (aem9296):

no you add it to the -24/5 to get -21/5

OpenStudy (anonymous):

Thats what I met, my bad. Thank you

OpenStudy (anonymous):

And I'm multiplying the denominator of the fraction to get it in standard form?

OpenStudy (aem9296):

yes standard form has to be whole numbers